MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2006/06/29
OBSERVER: Y. Matsubara T. Okumura F. Abe
CONDITION: fine
COMMENT: We changed valves for the vacuum on June 29th. This is to make the vacuum sensor indifferent to the camera when we want to replace a wrong sensor by a good one. The replacement finished at 4:00 pm. We started pumping after 4pm, and stopped the pumping at 1:am on June 30. The value of the vacuum wher we stopped observing at 7:30am, was 1.2E-3.
Compressor gas pressure: MPa
Expected Observation Time: START 18:23 END 07:19
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 1729-1738 | 10 | 60 | 3600 |
LMC | 533-534 | 2 | 8 | 2400 |
SMC | 137-137 | 1 | 9 | 2700 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 13 | 77 | 8700 (02h25m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
2 | 2 | 2 | 2 | 10 | 2 | 2 | 2 | 10 | 2 | 2 | 2 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
2 | 2 | 2 | 2 | 2 | 2 | 2 | 2 | 2 | 2 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 2 | 2 | 2 | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 |
0 | 0 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CHECK SHEET
Time | Status | Weather |
21:10 | waiting | cloudy |
0:25 | waiting | slightly cloudy |
3:20 | observing | fine |
4:10 | observing | fine |
5:06 | observing | slightly cloudy |
6:05 | observing | fine |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day